Problem 14 44ゆ Intro Apple stock had a return of 11%two years ago, and 22% last year. Google stock had a return of 12% two years ago, and 13% last year. Attempt 1/10 for 10 pts Part 1 Ifyou invest 30% in Apple and 70% in Google, what is the expected return of your porttolo for next year, if you use the past as a guide for the future (which is generally not a good idea for calculating expected returns)? Submit Attempt 1/10 for 10 pts Part 2 What is the expected variance of returns on Apple stock based on the historical data? 4+ decima Submit Attempt 1/10 for 10 pts. Part 3 What is the expected variance of returns on Google stock based on the historical data? 6+ decima Submit Attempt 1/10 for 10 pts Part 4 What is the expected covariance of returns based on the historical data? 5+ decima Submit Attempt 1/10 for 10 pts. Part 5 What is the correlation
Attempt 1/10 for 10 pts. Part 6 What is the variance of the risky portfolio? 5+ decima Submit Attempt 1/10 for 8 pts. Part 7 What is the standard deviation of the risky portfolio? 4+ decima Submit
Answer #1

Apple expected returns=(11%+22%)/2=16.50%

Google expected returns=(12%+13%)/2=12.50%

1.
=30%*16.50%+70%*12.50%=13.70%

2.
=sqrt((11%-16.50%)^2+(22%-16.50%)^2)/sqrt(2-1)=7.78%

3.
=sqrt((12%-12.50%)^2+(13%-12.50%)^2)/sqrt(2-1)=0.71%

4.
=(11%-16.50%)*(12%-12.50%)+(22%-16.50%)*(13%-12.50%)=0.00055

5.
=0.00055/(7.78%*0.71%)=0.995691372